OpenStudy (anonymous):

If you borrow $12,000 for 30 months at 6.5% simple interest, what is the total amount you will have to repay?

jimthompson5910 (jim_thompson5910):

use the formula I = P*r*t to find out how much interest I you'll have to pay back. P = 12000 is the principal r = 0.065 is the interest rate as a decimal t = 30/12 is the time in years notice how I did not simply write t = 30. The t = 30 represents 30 years and not 30 months, so that's why I divided by 12.

jimthompson5910 (jim_thompson5910):

Once you calculate the value of I, you add that to P to get P+I that is the total amount you pay back

jimthompson5910 (jim_thompson5910):

Or you could do it all in one formula A = P + P*r*t either way works